OpenStudy (anonymous):

Please solve the equations: x^4+4x^3=x^2+16x+12 x^4-3x^3+4x^2-6x+4=0

OpenStudy (anonymous):

for: x^4-3x^3+4x^2-6x+4=0 x=1 and x=2 the other are complex

OpenStudy (anonymous):

yes, but how i should count that?

OpenStudy (anonymous):

do you know how to do synthetic division yet?

OpenStudy (anonymous):

rather yes, but in this equation i can't do this :(

OpenStudy (anonymous):

solutions to polynomials like this takes a lot of guessing or try to graph them

OpenStudy (anonymous):

for this one x^4-3x^3+4x^2-6x+4=0 the roots maybe are: 1,3,4,2,etc you have to try to do synthetic divisions here using these as trials

OpenStudy (anonymous):

ok, so the roots depens from " ^ " ? yes?

OpenStudy (anonymous):

like for this one, x^4+4x^3=x^2+16x+12 simplify it first to x^4+4x^3-x^2-16x-12=0, now the possible roots are 1,-1,2,-2,3,-3,4,-4 etc, now do the synthetic divisions here using these numbers

OpenStudy (anonymous):

ok, but how i should to know these numbers?

OpenStudy (anonymous):

x^4-3x^3+4x^2-6x+4=0 the coefficients of the polynomials are, 1,-3,4,-6 did you see them? so the possible roots are + and negative of them

OpenStudy (anonymous):

positive or negative multiples of them

OpenStudy (anonymous):

ok, i see :) thanks for your help :)

OpenStudy (anonymous):

ok yw, good luck now

OpenStudy (anonymous):

thank you very much :)

OpenStudy (anonymous):

by the way, do your synthetic divisions and check your answer, x^4+4x^3-x^2-16x-12=0 for this the roots are -3.-2,-1,2 ans...

OpenStudy (anonymous):

yes, i have the same anwers, thanks once again :)

OpenStudy (anonymous):